OpenStudy (anonymous):

plz help . how is an electromagnetic wave able to travel through vaccum? i m unable to get this believed????????????

OpenStudy (anonymous):

@UnkleRhaukus @Yahoo! @experimentX

OpenStudy (anonymous):

em waves are able to travel thru space because they have a quantum particle associated with them. It is the particle travelling thru space, although the particle is of course massless and weightless. All quantum particles, including matter, are just packets of energy, and it's the properties of an individual particle that determines how it will behave (ie, will it behave like matter or energy). So light is able to exist on its own for the same reason matter can. Compressional waves cannot; they are simply a travelling perturbation of matter. http://uk.answers.yahoo.com/question/index?qid=20080302142609AAnzDv4

OpenStudy (anonymous):

but any kind of energy dont have its identity without matter

OpenStudy (experimentx):

Electric and Magnetic field traverse through vacuum. EM waves have two mutually inductive magnetic and electric field as it's two component that travel in wave.

OpenStudy (unklerhaukus):

(virtual) photons

OpenStudy (anonymous):

but photon is massless so acoo to einstien eq e=mc^2 photon bear no energy

OpenStudy (unklerhaukus):

a photon has energy \[E=h\nu\] where \(\nu\) is the frequency

OpenStudy (ujjwal):

energy of photon is not determined by their mass.. They don't have rest mass but have dynamic mass. However, energy of photon is given by \[E=h\nu\]

OpenStudy (anonymous):

but what about E=0*c^2=0

OpenStudy (experimentx):

\[ E^2 = m_0^2c^4 + p^2c^2 \] photon do no have mass but has momentum.

OpenStudy (experimentx):

*rest mass

OpenStudy (anonymous):

momentum =velocity*mass=0*c=0

OpenStudy (anonymous):

so how photon bears momentum

OpenStudy (anonymous):

can i say photon is particle of mattter???????

OpenStudy (ujjwal):

momentum= mc also \[mc^2=\frac{hc}{\lambda}\]or, \[mc=\frac{h}{\lambda}\]which is momentum of photon and yeah photon is like a particle..

OpenStudy (experimentx):

use wave-particle duality ... photos are able to give momentum to electrons while they do not have mass. while electrons having mass are able to travel in wave like manner. same goes for heavier particle. \[ \lambda = {h \over p} \implies {p = {h \over \lambda }} \\ E^2 = p^2 c^2 = {h^2c^2 \over \lambda^2 }\] you would get the same result postulated by plank.

OpenStudy (anonymous):

thanks a lot @experimentX & @ujjwal

OpenStudy (experimentx):

no probs!! I'm lack quite a lot on this topic.

OpenStudy (anonymous):

now my real question arises which i will ask as next question plz provde helping hand

OpenStudy (anonymous):

that means photon is carrier of all em waves

OpenStudy (experimentx):

your EM waves travelling though space ... you don't need to go to QM. it can be explained classically.

OpenStudy (anonymous):

or only light

OpenStudy (experimentx):

you know magnetic field and Electric field travel through vacuum. this should be enough to justify that EM waves travel through vaccum.

OpenStudy (anonymous):

i think i should not have closed this question shall i reopen a lot more to ask from this

OpenStudy (anonymous):

what happens when photoelectric effect takes place.here photon carries em wave

OpenStudy (ujjwal):

Photon doesn't carry wave.. It carries packets of energy! And since it behaves as particle, it knocks off electrons which is seen in photoelectric effect.. You can consider photon to be a ball and electron to be another ball. The photon literally knocks of electron and makes it free from atom thus causing photoelectric effect!

OpenStudy (ujjwal):

We usually don't mix 'wave concept' and 'particle concept'! And when we do it.. things get confusing.. At least for me.. While studying photo-electric effect, you just need to deal with particle concept!